I need an explicit formula for the sequence : {2, 6, 12, 20} I have noticed that the pattern is +4, +6, +8 and so on, but I can't figure out how to write the formula.
Answer by ad_alta(240)

Use A(n)=n(n+1) [since 1(1+1)=2, 2(2+1)=6)...4(4+1)=20]